OpenStudy (anonymous):

solve the equation and explain your steps? 2 r – 8 + 5 r = 10 + r

OpenStudy (anonymous):

2r-8+5r=10+r (add like terms)=> 7r-8=10+r 6r=18 >>> 18/6 r=3 Make sense?

OpenStudy (anonymous):

\[2r-8+5r=10+r\\7r-8=10+r\\6r-8=10\\6r=18\\r=3\]1. Combine like terms, so 2r + 5r = 7r. 2. Subtract r from both sides, so 7r - r = 6r. 3. Add 8 to both sides to isolate 6r. 4. Divide both sides by 6 to find r.

OpenStudy (anonymous):

thank you guys :)